Solve using the quadratic formula. 3x² - 5x = 5 Select the correct choice below and fill in any answer boxes in your choice.Please show work I want to learn! Simplify answer, using radicals as needed. Use integers or fractions for any number in the expression. Use a comma to separate answers as needed
Answer #1

Solution:

For a quadratic function \small ax^{2}+bx+c=0 ,

roots are given as:

\small x = \frac{-b\pm \sqrt{b^{2}-4ac}}{2a}

\small 3x^{2}-5x=5

\small 3x^{2}-5x-5=0

\small x = \frac{-(-5)\pm \sqrt{(-5)^{2}-4\times 3\times (-5)}}{2\times 3}

\small x = \frac{5\pm \sqrt{25+60}}{2\times 3}

\small x = \frac{5\pm \sqrt{85}}{6}

\small x = \frac{5+ \sqrt{85}}{6}\textup{ and }x = \frac{5- \sqrt{85}}{6}